# TiTok: Transfer Token-level Knowledge via Contrastive Excess to Transplant LoRA

### v1

@@ Appendix E Comparison with Additional Baseline
Appendix F Synthetic Data Generation with 2× Pool and Top-$M$ Selection
@@
In this section, we provide further details regarding the number of synthetic data samples used. In generating synthetic data, we first provide five samples from the original training set as few-shot exemplars to guide the model toward producing outputs in the desired style and format. For each task, we initially create a synthetic pool containing twice the number of examples used in the source model’s training. This pool is then filtered using contrastive excess scores, after which we retain only the top $M$ samples, where $M$ equals the size of the source training set. To be specific, we set $M=250$ for BBH, $M=90$ for MMLU, and $M=200$ for LaMP tasks. This procedure ensures that the target model’s LoRA adapter is trained on a dataset comparable in scale to that of the source model, while selective filtering enhances the overall quality of the retained data.
Appendix G ROUGE-L Filtering for Diverse Synthetic Queries
We now proceed to provide the task lists for which we did not apply ROUGE-L filtering when generating diverse synthetic queries. In general, we use ROUGE-L filtering to encourage diversity in queries, but for the tasks listed in Table 7, we only applied simple deduplication. In the case of BBH, tasks such as boolean_expressions or temporal_sequences already follow highly restricted and repetitive patterns, making high ROUGE-L scores inevitable. For MMLU, the only tasks without ROUGE-L filtering are the history-related subjects. This is potentially because history questions often require long passages that overlap in vocabulary, phrasing, or factual references (e.g., recurring names, dates, or events). Applying strict ROUGE-L filtering in such cases would make it difficult to generate the required number of synthetic queries. For all remaining BBH and MMLU tasks, as well as all LaMP tasks, we applied a ROUGE-L threshold of 0.7 to encourage diversity while still preserving task fidelity. Table 7 provides the tasks for which ROUGE-L filtering is not applied.
@@ Appendix G ROUGE-L Filtering for Diverse Synthetic Queries
BBH bbh_boolean_expressions, bbh_date_understanding, bbh_disambiguation_qa, bbh_geometric_shapes, bbh_logical_deduction_three_objects, bbh_multistep_arithmetic_two, bbh_navigate, bbh_object_counting, bbh_penguins_in_a_table, bbh_reasoning_about_colored_objects, bbh_salient_translation_error_detection, bbh_snarks, bbh_temporal_sequences, bbh_tracking_shuffled_objects_three_objects, bbh_web_of_lies
@@
Appendix H Prompts for Synthetic Query Generation
@@
In this section, we present the prompts used for generating synthetic queries. Each prompt includes five examples, which correspond to the seed prompts taken from the original data. For BBH, which consists of multiple subtasks, we specify task-specific instructions and formatting rules so that the generated queries and labels are structured according to the expected format. Each BBH subtask is paired with its corresponding rules and output format to ensure consistency across the benchmark, and Table 8 illustrates the case of boolean_expressions as a representative example.
@@
In contrast, for MMLU a single unified prompt format is sufficient for most subjects, and the model reliably generates queries that follow the expected structure. The exception is the history-related tasks, namely, high_school_us_history, high_school_world_history, and high_school_european_history. For these tasks, it is observed that more specific prompting is required to obtain better generations. The general MMLU prompt is shown in Table 9, whereas the history-related tasks make use of the specialized prompt provided in Table 10.
@@
Finally, for LaMP tasks, we apply the same query template uniformly to each user. The detailed prompt templates for LaMP 4 and 5 are provided in Tables 11 and 12 respectively.
@@
Table 8: Synthetic query prompt for the BBH boolean_expressions task. Each BBH subtask requires task-specific instructions and formatting rules to ensure that generated queries and labels follow the expected format. We show the boolean_expressions task here as a representative example.

### v2

@@ Appendix E Comparison with Additional Baseline
[…]
Table 11: Main results with variance. Mean $\pm$ standard deviation over three random seeds on BBH, MMLU, News Headline, and Scholarly Title under four transfer settings. BBH and MMLU are evaluated by exact-match accuracy, while News Headline and Scholarly Title are evaluated by ROUGE-1/L. All evaluations are zero-shot. Best scores are in bold, while second highest are underlined.
@@
[…]
Table 12: TiTok performance across $k=10\%$$90\%$. “Vanilla” denotes performance of target vanilla model. Highlighted columns denote the selected universal $k\%$ reported in Table 1. Highest performance are in bold, while the second best results are underlined.
[…]
Mistral 7B $\rightarrow$ Mistral 7B BBH Acc 0.397 0.401 0.420 0.444 0.445 0.441 $\cellcolorkSel0.432$ 0.431 0.431
[…]
News Headline R-1 0.117 0.153 0.159 0.161 0.161 0.160 $\cellcolorkSel0.160$ 0.160 0.161
R-L 0.101 0.138 0.143 0.144 0.143 0.142 $\cellcolorkSel0.142$ 0.142 0.142
Scholarly Title R-1 0.381 0.466 0.475 0.481 0.474 0.473 $\cellcolorkSel0.473$ 0.473 0.470
R-L 0.311 0.408 0.419 0.424 0.416 0.414 $\cellcolorkSel0.414$ 0.414 0.411
[…]
MMLU Acc 0.469 0.487 0.483 0.500 0.492 0.492 $\cellcolorkSel0.488$ 0.494 0.500
News Headline R-1 0.125 0.140 0.141 0.142 0.142 0.141 $\cellcolorkSel0.140$ 0.138 0.138
[…]
Scholarly Title R-1 0.444 0.460 0.458 0.458 0.460 0.467 $6 \cellcolorkSel0.4$64 0.465 0.465
[…]
MMLU Acc 0.469 0.479 $7 \cellcolorkSel0.4$75 0.475 0.475 0.474 0.475 0.474 0.472
News Headline R-1 0.125 0.127 $\cellcolorkSel0.127$ 0.128 0.122 0.121 0.122 0.121 0.121
[…]
Llama2 7B $\rightarrow$ Llama3 8B BBH Acc 0.469 0.527 0.519 0.531 $0.514 \cellcolorkSel0.510$ 0.509 0.509
[…]
R-L 0.110 0.110 0.114 0.115 0.119 $120 \cellcolorkSel0.1$22 0.121 0.121
[…]
Appendix F Extended Table 1 Results with Variance and Statistical Detail
We additionally report the mean ± standard deviation over 3 random seeds of our results in Table 1. The results are presented in Table 11. As shown, the overall trend remains consistent: relative to the target vanilla baseline, KD improves by 8.5%, TransLoRA by 4.44%, and TiTok by 9.94%. In fact, these improvements remain substantially larger than the corresponding standard deviations
Appendix G Comprehensive k% Sensitivity Analysis
In this section, we conduct an extensive experiment on TiTok’s performance throughout a wide range of k values, from 10% to 90%. The results are comprehensively presented in Table 12. These experiments demonstrate that TiTok consistently outperforms the target vanilla baseline throughout a broad, steady range of values, while the exact performance varies slightly depending on k%. This clearly indicates that the method is not overly sensitive to the choice of k% and that there exists a broad range of reasonable k% values where improvements are reliably obtained.
The reason for choosing a universal k% hyperparameter in our study is to keep the presentation of our paper more coherent and consistent. For this reason, we intentionally avoid task-specific tuning and report the k% that generally works well across tasks. However, we note that with additional per-task optimization of k%, further performance improvements are indeed possible.
With regard to the adaptive mechanism, each transfer setting exhibits its own effective k% range. For instance, same-backbone BBH transfer favors a mid-range k% (30–70%), weak-to-strong BBH transfer benefits from smaller k%, and weak-to-strong News Headline Generation shows the opposite trend, preferring larger k%. Taken together, these trends suggest that a simple adaptive mechanism can already be practical and effective. In our experiments, even coarse adjustments (i.e., using around 30% for weak-to-strong reasoning transfers and 70% for weak-to-strong stylistic tasks as denoted in Section 4.3) prove sufficient, without requiring extensive hyperparameter searches.
Appendix H Synthetic Data Generation with 2× Pool and Top-M Selection
@@
In this section, we provide further details regarding the number of synthetic data samples used. In generating synthetic data, we first provide five samples from the original training set as few-shot exemplars to guide the model toward producing outputs in the desired style and format. For each task, we initially create a synthetic pool containing twice the number of examples used in the source model’s training. This pool is then filtered using token-wise contrastive excess scores, after which we retain only the top $M$ samples, where $M$ equals the size of the source training set. To be specific, we set $M=225$ for BBH, $M=90$ for MMLU, and $M=200$ for LaMP tasks. This procedure ensures that the target model’s LoRA adapter is trained on a dataset comparable in scale to that of the source model, while selective filtering enhances the overall quality of the retained data.
Appendix I ROUGE-L Filtering for Diverse Synthetic Queries
[…]
@@ Appendix I ROUGE-L Filtering for Diverse Synthetic Queries
BBH boolean_expressions, date_understanding, disambiguation_qa, geometric_shapes, logical_deduction_three_objects, multistep_arithmetic_two, navigate, object_counting, penguins_in_a_table, reasoning_about_colored_objects, salient_translation_error_detection, snarks, temporal_sequences, tracking_shuffled_objects_three_objects, web_of_lies
@@
Appendix J Robustness of TiTok to Synthetic Data Quality
While TiTok’s primary contribution is token-level transfer and is thus not tied to synthetic data, we provide a detailed analysis on the robustness of our method to synthetic data variations.
Robustness to low-quality synthetic data. First, we deliberately construct a suboptimal synthetic dataset by selecting only the 250 lowest-scoring synthetic samples, as evaluated by the GPT-4o-mini grading model and prompt in (scores $\in\{0,1,2,3,4,5\}$) (Chen et al., 2024a). All training hyperparameters and filtering processes are identical to the original setup. As shown in Table 14, TiTok still substantially outperforms all baselines in the Mistral 7B → Mistral 7B BBH transfer setting. This shows that TiTok is still effective even with low-quality synthetic data.
Robustness to data diversity. Our synthetic data generation pipeline already incorporates deduplication and a ROUGE-L diversity threshold. This ensures the exclusion of exact duplicates and highly similar samples. More details are provided in Appendix I.
Robustness to prompt design. We further conduct an experiment with a different prompt for generating synthetic data. The alternative prompt template is provided in Table 15. We test this on BBH tasks in the Mistral 7B → Mistral 7B and Mistral 7B → Llama3 8B transfer settings.
Table 16 shows the results of using the alternative prompt for generating synthetic data. Analytically, the results show that TiTok achieves consistent improvements, and there are no substantial differences across prompt choices. This demonstrates that TiTok is not sensitive to prompt formulation, as long as the generated examples adhere to a reasonable and coherent structural pattern.
Overall, we emphasize that a highly refined or meticulously selected synthetic dataset is not required for TiTok. As long as the generated samples retain a relatively cohesive task structure, TiTok is still effective and is not very sensitive to the particular features of the synthetic data.
Model Transfer Method BBH (Acc.)
Mistral 7B $\rightarrow$ Mistral 7B Vanilla 0.397
KD 0.406
TransLoRA 0.405
TiTok (k=70%) 0.416
[…]
black
Table 14: BBH exact-match performance for the Mistral 7B $\rightarrow$ Mistral 7B transfer setting using low-quality synthetic data. TiTok (k=70%) continues to outperform all baselines.
Table 15: Alternative synthetic query prompt template for BBH.
Synthetic query generation prompt for {task_name} in BBH.
Generate {task_name} questions like these examples:
Example 1:
(few-shot example 1)
Example 2:
(few-shot example 2)
Example 3:
(few-shot example 3)
Example 4:
(few-shot example 4)
Example 5:
(few-shot example 5)
Example 6:
[…]
black
Table 16: Effect of prompt variation on TiTok on BBH in the Mistral 7B → Mistral 7B and Mistral 7B → Llama3 8B settings.
Transfer Method Prompt BBH (Acc.)
Mistral 7B $\rightarrow$ Mistral 7B Vanilla 0.397
TiTok (k=70%) original (Table 22) 0.432
alternative (Table 15) 0.436
Mistral 7B $\rightarrow$ Llama3 8B Vanilla 0.469
TiTok (k=70%) original (Table 22) 0.482
alternative(Table 15) 0.485
[…]
black
Appendix K Addressing Synthetic Data Quality and Bias
We now move on to explain how our pipeline incorporates several components designed to reduce bias and prevent quality degradation in synthetic data. Specifically, we achieve this by incorporating several components, including a length filter that removes malformed or low-informative samples, a token-wise contrastive excess mechanism that selects only the most informative tokens rather than relying on the entire synthetic instances, and diversity filtering ( i.e., deduplication and ROUGE-L thresholding) to prevent mode collapse during synthetic data generation.
Additionally, we further execute an external quality evaluation using a separate model (GPT-4o-mini), adopting a robust synthetic data assessment methodology established in a prior work (Chen et al., 2024a). Table 17 reports the average quality scores (scale 1–5) of gold data and the synthetic data generated by multiple source models used in our experiments.
Data Source BBH Scholarly Title
Gold data 2.40 3.68
Mistral 7B synthetic data 2.27 3.89
Llama3 3B synthetic data 2.01 3.74
Llama2 7B_synthetic data 2.07 3.15
[…]
black
Table 17: Overall average quality ratings of gold and synthetic datasets across the BBH and Scholarly Title Generation tasks. GPT-4o-mini evaluates each data sample on a 1–5 scale.
These results show that the synthetic data used in our experiments is comparable in quality to the gold dataset, with no evidence of harmful bias dominating the data. Importantly, TiTok performs consistently well across all synthetic sources, even when the quality varies (e.g., 2.01 vs. 2.40 in BBH), which further indicates that TiTok does not rely on perfectly clean or unbiased synthetic data. Taken together, these findings demonstrate that TiTok is robust to imperfections in synthetic data, and, once again, can be effectively used even in non-synthetic datasets as shown in Table 4.
Appendix L Exploration on Self Refinement
While TiTok primarily focuses on LoRA-to-LoRA knowledge transfer, we investigate its potential for self-refinement. We perform this experiment on BBH in the Mistral 7B → Mistral 7B and the Mistral 7B → Llama3 8B transfer settings. In particular, after the initial transfer step, 1) we generate synthetic data using the target LoRA that had already learned TiTok’s transferred knowledge, 2) compute token-wise contrastive excess scores on the newly generated synthetic data using the original source LoRA and the its base model, and 3) train a new target LoRA with the same hyperparameters in the original setting (i.e., k=0.7%). The results are summarized in Table 18.
Table 18: “Initial TiTok” indicates TiTok applied once. “Iterative TiTok: self transfer” applies TiTok again using synthetic data generated by the learned target LoRA after the initial transfer.
BBH (Acc.)
Mistral 7B $\rightarrow$ Mistral 7B Vanilla 0.397
Initial TiTok 0.432
Iterative TiTok: self transfer 0.456
Target expert 0.460
Mistral 7B $\rightarrow$ Llama3 8B Vanilla 0.469
Initial TiTok 0.482
Iterative TiTok: self transfer 0.510
Target expert 0.531
Interestingly, the results show that the iterative step yields additional improvement, implying that iterative refining is possible. We hypothesize that this is possible because the initial transfer increases the target model’s familiarity with the task and partially aligns its internal representations with the source domain. As a result, the target model generates cleaner, more coherent synthetic data with fewer irrelevant patterns. This thereby enables the second iteration of TiTok to more accurately identify domain-informative tokens, which improves token selection and makes excess scoring more discriminative. However, we emphasize that iterative refining synthetic data cannot surpass a target model trained on real train data. By creating increasingly better synthetic data, iteration can help narrow the gap, but it is still far from a fully supervised target model trained on actual data.
Appendix M Robustness of TiTok Under Weaker Source LoRAs
We now discuss scenarios where the source LoRA is weaker than the target model. While it is true that TiTok uses the source LoRA as a reference, its effectiveness is not strongly dependent on the source LoRA’s absolute performance. To illustrate this point more clearly, Table 19 summarizes the subsets of Table 1 in which the source LoRA underperforms the target model’s vanilla version, and presents these source LoRA scores alongside the corresponding transfer results.
As shown in the table, TiTok still consistently improves the target model even when the source LoRA is weaker than the target baseline. This shows that TiTok captures domain-specific signals that are embedded in key tokens rather than inheriting the source LoRA’s full behavior.
BBH Scholarly_Title
Model Transfer Metric Acc. R-1 R-L
Llama3 3b $\rightarrow$ Llama3 8b source lora 0.460 0.429 0.363
target base 0.462 0.444 0.378
TiTok (k=30%) 0.509 0.457 0.392
Llama2 7B $\rightarrow$ Llama3 8B source lora 0.359 0.431 0.367
target base 0.469 0.444 0.378
TiTok (k=70%) 0.510 0.461 0.404
[…]
Table 19: Results on BBH and Scholarly Title Generation for transfer settings where the source LoRA is weaker than the target baseline.
KD-based approaches, however, are highly dependent on the absolute capacity of the source LoRA. By making the target model follow the teacher’s logits, a weak or poorly tuned source LoRA unavoidably transfers certain flaws, resulting in the minimal gains supported by the results in the above table. Meanwhile, TiTok avoids these deficiencies because it fundamentally differs in how it leverages the source LoRA. Instead of relying on the source LoRA’s logits globally, TiTok uses the source only to smartly identify informative token positions. As a result, a poorly tuned source LoRA does not mislead TiTok since it does not make the target model imitate the source LoRA’s behavior. Rather, it selectively retrieves the useful tokens using the source LoRA’s domain knowledge.
Appendix N Computational Overhead Comparison
Here, we present the computational differences between TiTok and TransLoRA. We conduct an end-to-end time comparison of TransLoRA and TiTok on the BBH and News Headline Generation benchmarks in the Mistral 7B → Llama3 8B transfer setting. The results are presented in Table 20 and 21. Across both benchmarks, TiTok achieves roughly a 1.5×–2.5× reduction averagely in total compute time. Most of the gain comes from the removal of discriminator training, while the per-token log-likelihood computation produces only a minimal and manageable overhead.
BBH (250 data × 27 tasks)
[…]
Table 20: Runtime comparison for Mistral 7B $\rightarrow$ Llama3 8B transfer on BBH.
[…]
black
News_Headline (200 data × 30 users)
[…]
Table 21: Runtime comparison for Mistral 7B $\rightarrow$ Llama3 8B transfer on News Headline Generation.
Appendix O Qualitative Examples
In this section, we provide qualitative examples from the News Headline Generation in the Mistral 7B → Mistral 7B transfer setting to illustrate how TiTok identifies the most informative tokens. The qualitative examples are presented in Figure 5. Analytically, the selected tokens correspond to the beginnings of major semantic or structural units in the headline. These tokens carry the highest information value: they denote transitions, introduce core noun phrase components, or contain root words. Moreover, since the token-wise contrastive excess score is defined as the difference between the expert model (source LoRA + base) and the amateur model (base only), it pinpoints the tokens where the expert and amateur diverge most in their predictions and so include effective learning signals for target tasks. Taken together, these qualitative examples effectively demonstrate that TiTok consistently retains the tokens that are most important for providing structure and meaning.
Figure 5: Qualitatave examples from News Headline Generation in the Mistral 7B → Mistral 7B setting. Bold and highlighted tokens are part of the selected tokens. Notably, these tokens exhibit high token-wise contrastive excess scores and thus fall into the top 70% selected for training.
Appendix P Prompts for Synthetic Query Generation
@@
In this section, we present the prompts used for generating synthetic queries. Each prompt includes five examples, which correspond to the seed prompts taken from the original data. For BBH, which consists of multiple subtasks, we specify task-specific instructions and formatting rules so that the generated queries and labels are structured according to the expected format. For example, Table 22 illustrates the case of boolean_expressions as a representative example in detail.
@@
In contrast, for MMLU, a single unified prompt format is sufficient for most subjects. The exceptions are the history-related tasks, namely, high_school_us_history, high_school_world_history, and high_school_european_history. For these tasks, it is observed that more specific prompting is required to obtain better generations. The general MMLU prompt is shown in Table 23, whereas the history-related tasks make use of the specialized prompt provided in Table 24.
@@
Finally, for LaMP tasks, we apply the same query template uniformly to each user. The detailed prompt templates used for this benchmark are provided in Tables 25 and 26.
[…]
black
@@
Table 22: Synthetic query prompt for the BBH boolean_expressions task. Each BBH subtask requires task-specific instructions and formatting rules to ensure that generated queries and labels follow the expected format. We show boolean_expressions task here as a representative example.